OpenStudy (anonymous):

1. Solve the system of equations. y = 2x² – 3 y = 3x – 1

hero (hero):

Hint: set y = y

OpenStudy (anonymous):

ans is 2 and -1/2

hero (hero):

You might want to include the steps along with that

OpenStudy (anonymous):

i don't know how to add exponents on a number i am trying can u tell

hero (hero):

No, I cannot tell because all you have posted is an answer. If that was the case, I already had that answer two seconds after the question was posted.

hero (hero):

You're supposed to give guidance, explanations, and steps, not answers.

OpenStudy (anonymous):

i know dude, i am trying to type the explanation. first time i entered this site. so i don't know how to add exponents like 2x now i don't know how to add power 2 to the x. i am trying this since the question was on the site

hero (hero):

Well, I have the steps. I will post them

OpenStudy (anonymous):

now u can see details

hero (hero):

\[y = y\]\[2x^2 - 3 = 3x - 1\]\[2x^2 - 3x - 3 + 1 = 0\]\[2x^2 - 3x - 2 = 0\]\[2(x^2 - \frac{3x}{2} - 1) = 0\]\[x^2 - \frac{3x}{2}-1 = 0\]\[x^2 - \frac{3x}{2} = 1\]\[x^2 - \frac{3x}{2} + \frac{9}{16} = 1 + \frac{9}{16}\]\[(x - \frac{3}{4})^2 = \frac{16}{16} + \frac{9}{16}\]\[x - \frac{3}{4} = \sqrt{\frac{25}{16}}\]\[x = \pm \frac{5}{4} + \frac{3}{4}\]\[x = \frac{8}{4} = 2\]\[x = \frac{3}{4} - \frac{5}{4} = -\frac{2}{4} = -\frac{1}{2}\]

OpenStudy (anonymous):

very long process see my attachment.

hero (hero):

Yes, but it looks better

OpenStudy (anonymous):

u r beautifying the solution

hero (hero):

That's what I'm supposed to do. Mathematical solutions are supposed to look beautiful.

hero (hero):

lol

OpenStudy (anonymous):

\[2x ^{2}-3=3x-1\] \[2x ^{2}-3x-3+1=0\] \[2x ^{2}-3x-2=0\] \[2x ^{2}-(4x-x)-2=0\] \[2x ^{2}-4x+x-2=0\] \[2x(x-2)+1(x-2)=0\] \[(2x+1)(x-2)=0\] (x-2)=0, x=2 2x+1=0, x= -1/2

OpenStudy (anonymous):

this is the simple way to solve the problem. maths is not to beautify it is to understand the things and concept. you should try to solve the problems in easy manner so that everyone will get the good thing, if not then u should learn the things. u should not fight the people that u r doing the right thing. if u do so u never learn the new things.

hero (hero):

Look how long it took you to post the solution. If I had to do it all over again, I know which method I would use.

hero (hero):

The truth is, my method is faster than your method. Plus mine doesn't involve guessing.

OpenStudy (anonymous):

i know u know about urself. it looks the time because i was unaware how to type this. but when i got it i typed the solution. learning different methods will not harm u. and adopting the methods which is easy will also not harm you. well its on you... people will decide this whose method is better.

hero (hero):

I already made the decision

OpenStudy (anonymous):

good...